# Search and Society Reimagining Information Access for Radical Futures

### v1

@@ 4 Towards sociotechnical change
Reflections for IR. The survey of works presented in this section hopefully provides some seeds of ideas for how IR research can be driven by radical new sociotechnical imaginaries. This is not to imply that these other IR-adjacent fields have achieved the desired success from these approaches, in fact there are some evidence [76] that point otherwise. Rather, we should recognize that how values and politics can inform computing research is still an open question, and they may apply differently to IR than these other fields. The challenge then for the community is to collectively engage and push towards sociotechnical change. In the remaining of this section, we discuss how we imagine some of these frames and values can guide us towards open challenges in information access. However, these examples should be interpreted as just that, as examples, not our recommendation for specific research questions the community should focus on.
@@
Through the lens of feminist, queer, and anti-racist IR, we could critique existing approaches to ranking fairness, not only in terms their use of socially constructed categories, such as race and gender [288], but question if it is the appropriate framing at all for the problems it purports to solve. For example, instead of trying to algorithmically fix under-representation of women and people of color in image search results for occupational roles, we could reclaim that digital space as a site of resistance and emancipatory pedagogy by allowing feminist, queer, and anti-racist scholars, activists, and artists to create experiences that teach the history of these movements and struggles.
@@
Anti-oppressive IR research may concern itself with questions such as: Can we translate Freire’s [133] anti-oppressive pedagogy to strategies for anti-oppressive information access? Can search result pages support dialogical interactions between searchers that allows for communities of searchers to add context to the search results, as an alternative to centralized moderation? Unlike conversation search, that is framed as interaction between the user and the system, the idea of dialogical search interfaces challenges us to build sophisticated sociotechnical solutions to support dialog between searchers in context of specific search intents in ways that leads to knowledge production and better digital literacy. Anti-oppressive and anti-capitalist perspectives may also motivate us to reimagine search and recommender systems as decentralized and federated [322].
@@
IR research may also employ these lenses as instruments of critique. For example, in the enterprise context, Gausen et al. [143] adopt decolonial and anti-capitalist lens to expose how information and knowledge access systems may commodify and appropriate knowledge from workers. We should also critically challenge the employment of what Gray and Suri [159] calls Ghost Work in IR research both as a labor issue and through the lens of decolonization. In abolitionist IR, we must ensure that the techonologies we build cannot be used for surveillance or any other military or carceral applications.
Finally, the community may also consider more radical direct actions such as developing critical theories of information access, or collectively organizing to abolish big tech [219].
5 How do we get there?
@@
In both academic research and industrial deployment, IR places a strong emphasis on the needs of the user (consumer). This focus motivates various lines of research including: understanding user needs (through lab studies, log analysis, surveys, etc.), optimizing the search system towards those needs (e.g., relevance optimization, personalization, and improving response time), and validating that proposed system changes indeed benefit the user (through online experimentations and further lab studies). Salient in industry settings are the needs of the system owners—e.g., revenue, market share, and brand—that drive significant decisions for system design and deployment, but have historically been of lesser concern to academic IR research. Real-world IR system deployments also engage with content producers and publishers—e.g., web publishers and the search engine optimization community for web search engines, and artists for music recommender systems), although how their needs are weighed against the needs of system owners and users may vary (e.g., [164, 290]). IR research has considered questions of fairness between producers, but have rarely focused on the power differentials between system owners and producers, and its implications for producers.
Societal concerns in both IR research and industry settings have commonly been framed through a narrow lens of harm mitigation, such as “how do we make the IR system more fair?” and “how do we reduce misinformation in search results?”, without fundamentally challenging the frames in which these systems are designed and deployed—e.g., centralized control and profit incentives [241, 344]. IR systems are deeply embedded in sociopolitical and organization context. However, instead of grounding IR research in questions around its role in online and institutional knowledge production, literacy and informed citizenry, public health education, and social justice, the community has typically constrained themselves to improving measurable system attributes like relevance and efficiency.
@@
Figure 1: Hierarchy of IR stakeholder needs. More fundamental and critical needs are at the bottom of the pyramid. This figure is inspired by Maslow’s Hierarchy of Needs [247] and Siksika (Blackfoot) way of life [303].
@@
– Lilla Watson and other members of an Aboriginal Rights group in Queensland [^8] While often credited for this quote, Watson explained [304] that this came out of a collective process by an Aboriginal rights group in Queensland that she was part of.
Our call for re-centering societal needs in IR research must not be confused for a call for technosolutionism. Quite the contrary, we believe we need safe spaces where IR researchers can engage with social scientists, legal scholars, critical theorists, activists, and artists, in a recognition of a collective struggle to develop shared understandings of the core challenges and what IR research can offer to this process. We must approach these spaces with curiosity and humility; in recognition of our own incomplete understanding of the world; open to change and be changed by these encounters.
@@
The mushroom at the end of the world: On the possibility of life in capitalist ruins [353]
Through these exchanges we must build relations of solidarity; work together to articulate pluralistic desirable sociotechnical futures, co-develop theories of change and new research agendas to support our aspirations; vigilantly critique our own assumptions and the structures that we exist in and conduct our research in; and critically assess the impact of our work not by publications or scholarly metrics but in terms of affecting real social change. This praxis will be unfamiliar and the learning curve frustrating; but hopefully the mutual shaping of IR techonology and society ultimately rewarding.
6 Who should do this work?
In his highly influential work “Pedagogy of the oppressed” [133], Freire criticizes the “banking” concept of eduction: “Education thus becomes an act of depositing, in which the students are the depositories and the teacher is the depositor. […] knowledge is a gift bestowed by those who consider themselves knowledgeable upon those whom they consider to know nothing. […] But the humanist, revolutionary educator cannot wait for this possibility to materialize. From the outset, her efforts must coincide with those of the students to engage in critical thinking and the quest for mutual humanization. His efforts must be imbued with a profound trust in people and their creative power. To achieve this, they must be partners of the students in their relations with them.” These profound words present a critique that we believe is also relevant to how we conduct fairness and ethics research in the IR community today.
@@ 6 Who should do this work?
To challenge the homogeneity of the future imaginaries—saliently bound by colonial, cisheteropatriarchal, and capitalist ways of knowing the world—that shape our research, we need broad and diverse participation from our community. But it is also in that very context that we must critically reflect on the topic of membership in our community itself. ACM SIGIR has a commendable emphasis on the topic of Diversity, Equity, and Inclusivity (DEI) [362, 210, 149, 151, 150]. For our sociotechnical imaginaries to be informed by pluralistic social, cultural, and political perspectives we not only need significantly improved representation from historically underrepresented and marginalized groups in our community, but also be inclusive of their politics and world views. Inclusion of people without inclusion of their history and struggles is simply tokenism and epistemic injustice [134]. That is why we believe that we, as a community, should go beyond just diversity and inclusivity (D&I), and enshrine as our goal Justice, Equity, and Diversity & Inclusivity (JEDI)—in which context D&I is both a means towards justice and equity, and also an end in itself. As Keyes et al. [208] put it: “This must be about more than just bodies: it is not diversity if we only accept marginalised people who are stripped of the epistemic models that underpin experiences of being Other, or have the work they draw from those models held to an unequal standard of legitimacy”.
@@
Lastly, we reiterate the important role of industry researchers in this process. They should take advantage of their proximity to large-scale systems to identify, understand, and communicate concerns of societal import and partner with academia to work on those challenges. The spaces they occupy are also sites for resistance [380, 381].
7 Discussion and conclusion

### v4

@@ 3.2 Proposals for IR
The survey of works presented in Section 3.1 hopefully provides some seeds of ideas for how IR research can be driven by radical new sociotechnical imaginaries. This is not to imply that these other IR-adjacent fields have achieved the desired success from these approaches, in fact there are some evidence (Chivukula and Gray, 2020) that point otherwise. Rather, we should recognize that how values and politics can inform computing research is still an open question, and they may apply differently to IR than these other fields. The challenge then for the community is to collectively engage and push towards sociotechnical change. In the remaining of this section, we discuss how we imagine some of these frames and values can guide us towards open challenges in information access. However, these examples should be interpreted as just that, as examples, not our recommendation for specific research questions the community should focus on. The actual research agenda should be developed through participatory processes that simultaneously focuses on both identifying technical research questions and building diverse communities with shared understanding of these challenges and shared commitments to address them.
@@
Through the lens of feminist, queer, and anti-racist IR, we could critique existing approaches to ranking fairness, not only in terms their use of socially constructed categories, such as race and gender (Pinney et al., 2023), but question if it is the appropriate framing at all for the problems it purports to solve. For example, instead of trying to algorithmically fix under-representation of women and people of color in image search results for occupational roles, we could reclaim that digital space as a site of resistance and emancipatory pedagogy by allowing feminist, queer, and anti-racist scholars, activists, and artists to create experiences that teach the history of these movements and struggles.
@@
Anti-oppressive IR research may concern itself with questions such as: Can we translate Freire’s (Freire, 2020) anti-oppressive pedagogy to strategies for anti-oppressive information access? Can search result pages support dialogical interactions between searchers that allows for communities of searchers to add context to the search results, as an alternative to centralized moderation? Unlike conversation search, that is framed as interaction between the user and the system, the idea of dialogical search interfaces challenges us to build sophisticated sociotechnical solutions to support dialog between searchers in context of specific search intents in ways that leads to knowledge production and better digital literacy. Anti-oppressive and anti-capitalist perspectives may also motivate us to reimagine search and recommender systems as decentralized and federated.
@@
IR research may also employ these lenses as instruments of critique. For example, in the enterprise context, Gausen et al. (2023) adopt decolonial and anti-capitalist lens to expose how information and knowledge access systems may commodify and appropriate knowledge from workers. We should also critically challenge the employment of what Gray and Suri (2019) calls Ghost Work in IR research both as a labor issue and through the lens of decolonization. In abolitionist IR, we must ensure that the techonologies we build cannot be used for surveillance or any other military or carceral applications. The community may also consider more radical direct actions such as developing critical theories of information access, or collectively organizing to abolish Big Tech (Kwet, 2020).
3.3 Where do we start?
We are calling for not only a significant shift in what the IR community works on, but fundamentally changing the arrangements within our community that determine on an ongoing basis our research agendas. Beyond explicating our values and sociotechnical imaginaries, we need to develop frameworks that help us appropriately prioritize societal needs against the needs of the user, the publisher, and the platform owners. We also need new research that reimagines how IR can be informed by different epistemologies and political theories. Finally, we must also critically reexamine the arrangements within our community and create spaces for shared sense-making in collaboration with those outside of our field. We elaborate on these further in this section.
@@ 3.3 Where do we start?
In both academic research and industrial deployment, IR places a strong emphasis on the needs of the user (consumer). This focus motivates various lines of research including: understanding user needs (through lab studies, log analysis, surveys, etc.), optimizing the search system towards those needs (e.g., relevance optimization, personalization, and improving response time), and validating that proposed system changes indeed benefit the user (through online experimentations and further lab studies). Salient in industry settings are the needs of the system owners—e.g., revenue, market share, and brand—that drive significant decisions for system design and deployment, but have historically been of lesser concern to academic IR research. Real-world IR system deployments also engage with content producers and publishers, e.g., web publishers and the search engine optimization community for web search engines, and artists for music recommender systems); although how their needs are weighed against the needs of system owners and users may vary, e.g., (Guttenberg, 2012; Plaugic, 2015). IR research has considered questions of fairness between producers, but have rarely focused on the power differentials between system owners and producers, and its implications for producers.
@@
Figure 1: Hierarchy of IR stakeholder needs. More fundamental and critical needs are at the bottom of the pyramid. This figure is inspired by Maslow’s Hierarchy of Needs (Maslow, 1958) and Siksika (Blackfoot) way of life (Ravilochan, 2021).
Societal concerns in both IR research and industry settings have commonly been framed through a narrow lens of harm mitigation, such as “how do we make the IR system more fair?” and “how do we reduce misinformation in search results?”, without fundamentally challenging the frames in which these systems are designed and deployed, e.g., centralized control and profit incentives (Mager, 2012; Taplin, 2017). IR systems are deeply embedded in sociopolitical and organization context. However, instead of grounding IR research in questions around its role in online and institutional knowledge production, literacy and informed citizenry, public health education, and social justice, the community has typically constrained themselves to improving measurable system attributes like relevance and efficiency.
@@
To realize structural changes in IR and meaningfully challenge dominant imaginaries we must also invest in research that specifically explores and builds on the connections between IR and different epistemologies and social and political theories. Examples of such work may consider how IR system design can support different models of democracy (Vrijenhoek et al., 2021; Helberger, 2021) and emancipatory aspirations (Mitra, 2025). We must also ensure that such research does not happen in isolation but is grounded in a collective effort to build a movement within the IR community. This requires us to acknowledge and value community building as an important part of IR research. In this context, it is important to emphasize that our call for re-centering societal needs in IR research must not be confused for a call for technosolutionism. Quite the contrary, we believe we need safe spaces where IR researchers can engage with scholars from other diverse fields, legal and policy experts, activists, and artists, in a recognition of a collective struggle to develop shared understandings of the core challenges and what IR research can offer to this process. These spaces are critical for cross-pollination of ideas and shared sense-making, which are vital for realizing structural changes. As Tsing aptly puts it:
@@
The mushroom at the end of the world:
On the possibility of life in capitalist ruins
 (Tsing, 2015)
Through these exchanges we must build relations of solidarity; work together to articulate pluralistic desirable sociotechnical futures, co-develop theories of change and new research agendas to support our aspirations; vigilantly critique our own assumptions and the structures that we exist in and conduct our research in; and critically assess the impact of our work not by publications or scholarly metrics but in terms of affecting real social change. In other words, before we can transform our platforms and systems we need to transform our communities and how we conduct research. And it is vital that we approach these spaces with curiosity and humility; in recognition of our own incomplete understanding of the world; open to change and be changed by these encounters. In this context, it would do IR researchers good to keep in mind the words of activist Lilla Watson.
– Lilla Watson and other members of an
Aboriginal Rights group in Queensland [^5] While often credited for this quote, Watson explained (Reblog of report from Northland Poster, ) that this came out of a collective process by an Aboriginal rights group in Queensland that she was part of.
This praxis will be unfamiliar and the learning curve frustrating; but hopefully the mutual shaping of IR techonology and society will be ultimately rewarding for us all. And we must embark on this transformative journey fully acknowledging that challenging dominant imaginaries is challenging power itself. To affect counter-imaginaries we must therefore actualize counter-structures and alternative funding mechanisms that can sustain this research in the face of likely reprisals from those whose power and visions are threatened by our proposed transformations.
3.4 Who should do this work?
In his highly influential work “Pedagogy of the oppressed”, Freire (2020) criticizes the “banking” concept of education: “Education thus becomes an act of depositing, in which the students are the depositories and the teacher is the depositor. […] knowledge is a gift bestowed by those who consider themselves knowledgeable upon those whom they consider to know nothing. […] But the humanist, revolutionary educator cannot wait for this possibility to materialize. From the outset, her efforts must coincide with those of the students to engage in critical thinking and the quest for mutual humanization. His efforts must be imbued with a profound trust in people and their creative power. To achieve this, they must be partners of the students in their relations with them.” These profound words present a critique that we believe is also relevant to how we conduct fairness and ethics research in the IR community today.
@@ 3.4 Who should do this work?
To challenge the homogeneity of the future imaginaries—saliently bound by colonial, cisheteropatriarchal, and capitalist ways of knowing the world—that shape our research, we need broad and diverse participation from our community. But it is also in that very context that we must critically reflect on the topic of membership in our community itself. ACM SIGIR has a commendable emphasis on the topic of Diversity, Equity, and Inclusivity (DEI) (Verberne et al., 2024; Kobayashi, 2017; Goharian and Bast, 2022; Goharian et al., 2023, 2021). For our sociotechnical imaginaries to be informed by pluralistic social, cultural, and political perspectives we not only need significantly improved representation from historically underrepresented and marginalized groups in our community, but also be inclusive of their politics and world views. Inclusion of people without inclusion of their history and struggles is simply tokenism and epistemic injustice (Fricker, 2007). That is why we believe that we, as a community, should go beyond just diversity and inclusivity (D&I), and enshrine as our goal Justice, Equity, and Diversity & Inclusivity (JEDI)—in which context D&I is both a means towards justice and equity, and also an end in itself. As Keyes et al. (2019) put it: “This must be about more than just bodies: it is not diversity if we only accept marginalised people who are stripped of the epistemic models that underpin experiences of being Other, or have the work they draw from those models held to an unequal standard of legitimacy”.
@@
Lastly, we reiterate the important role of industry researchers in this process. They should take advantage of their proximity to large-scale systems to identify, understand, and communicate concerns of societal import and partner with academia to work on those challenges. The spaces they occupy are also sites for resistance (Wickenden, 2018; Widder et al., 2023).
4 Why now?
The arguments we present in this paper to reimagine our sociotechnical futures and center IR research on societal needs have always been relevant to the field. However, there is a confluence of several factors that makes this discourse particularly relevant in the present moment. Research communities constantly evolve, shaped by ideas and developments from both within the field and adjacent communities, and in response to real world events and changing societal needs. In case of IR, we believe we are seeing significant developments in both context at present: (i) A fast-changing landscape across IR and adjacent fields, such as natural language processing (NLP), HCI, ML, and AI, spurred by recent progress in LLMs and other generative AI approaches, and (ii) an increasing recognition of the role of technology, and the communities that build it, in determining our collective futures. Consequently, there is a shared sense in the community that right now both IR technologies and IR research have been made malleable and are undergoing transformative changes under these forces of emerging new computational capabilities and evolving societal needs (Azzopardi et al., 2024a). This presents a timely opportunity for the field to consciously, collectively, and ambitiously engage in purposeful dialog about the future of the field while metaphorically the “IR(on)” is hot and before it is irrevocably shaped by unexamined imaginaries of those with power and influence over present day IR research. In doing so, we must also critically reflect on “where do we want to go?” (i.e., our sociotechnical imaginaries), “how do we get there?” (i.e., our theories of change), and “who will we go there with?” (i.e., our relationships, and that of our work, with other disciplines, governments, industry, and society). These considerations should drive future IR research as a whole, and we should accept this opportunity to re-center our research agendas on societal needs while dismantling the artificial separation between the work on fairness and ethics in IR and the rest of IR research.
LLMs are changing how we access information.
The natural language generation capabilities of LLMs are having a profound effect on how we access information and in what context. Conversational search interfaces have gone from being aspirational (Anand et al., 2021; Metzler et al., 2021) to being deployed at web-scale (e.g., Bing Chat [^6] https://chat.bing.com/ (now Microsoft Copilot (Mehdi, 2023)) and Google Bard [^7] https://bard.google.com/) in a span of two years. While the long-term social implications of inserting an LLM between a retrieval system and the information seeker should rightly be met with rigorous skepticism (Shah and Bender, 2022), natural language interfaces are already impacting how we interact with IR systems. The ubiquitous search box is being challenged as IR becomes more context-driven than user-driven as a consequence of LLMs increasingly embedding themselves in user’s work processes—e.g., Microsoft Copilot for M365 (Mehdi, 2024; Warren, 2024)—and interacting with the IR system on the user’s behalf, under retrieval-augmentation (Lewis et al., 2020; Zamani et al., 2022).
While we should be excited with the new prospects that these emerging AI technologies unlock and recognize that they will shape how we access and interact with information in the future, we must not be duped by AI techno-determinism into believing that there is a single pre-determined path forward. Instead, we must hold pluralistic views of what IR’s future, one that is yet to be determined, looks like and how these technologies will take us there. In a study of top-cited AI papers, many of which are coauthored by researchers affiliated with industry or elite universities, Birhane et al. (2022) find that the dominant values expressed and operationalized support concentration of power. So, we must ask: in what new ways can we imagine accessing and interacting with information, aided by LLMs, if large-scale IR systems were not just a purview of Big Tech? How would LLMs empower us to reimagine IR systems whose explicit goal is to dismantle hierarchies and redistribute power, not to centralize it? What role would AI technologies play in information access that is built explicitly to facilitate dialogical social processes for knowledge production, world building, and our collective struggles for universal emanicipation? It is critical that we have these conversations now in the face of ongoing massive technology-driven power shifts in favor of dominant established platforms that grants their visions of the future normative status and shrinks the space for any critique, resitence, or counter-imaginaries.
LLMs are shifting priorities of IR research.
Over the last decade, deep learning technologies became the new hammer in the toolbox for IR research (Mitra and Craswell, 2018; Lin et al., 2020; Fan et al., 2022), dominating IR publications with nearly four out of five papers at the ACM SIGIR 2020 conference being related to deep learning by some estimates (Mitra, 2021). One particular focus of neural IR has been on estimating relevance of information artefacts (e.g., documents) to an information intent (e.g., as expressed by a search query) for ranking, a central problem in IR. Curiously, many of the key ingredients for this research, such as the Transformer architecture (Vaswani et al., 2017) and the idea of pretrained LLMs, like BERT (Devlin et al., 2019), came from fields adjacent to IR; correspondingly, shifting the focus within the field more towards adapting these models for the relevance estimation task—e.g., (Nogueira and Cho, 2019)—and making them more efficient (Fröbe et al., 2024).
More recently, Thomas et al. (2024) demonstrated that LLMs, like GPT-4 (OpenAI, 2023), are able to estimate the actual searcher’s preference for documents, given their query, better than several populations of human relevance assessors. This technology has already been deployed in production at Bing. [^8] https://twitter.com/IR_oldie/status/1659413086007328768 Putting it bluntly, these LLMs may be getting close to the best we can expect with machine learned general purpose relevance estimators. If these claims stand the test of time, it may mark a watershed moment for IR research. Speculatively, we may see the IR community further shifting towards: (i) Improving efficiency of these models, (ii) focusing on more specialized information needs—e.g., tip-of-the-tongue information needs (Arguello et al., 2021), and (iii) increasing investments in measurement and evaluation—e.g., for emerging new IR scenarios, such as retrieval-enhanced machine learning (Zamani et al., 2022). Alternatively, we may ask: How can the IR community meet this moment, not with apprehension nor with unchallenged exuberance for progress happening in adjacent fields, but truly grasp this opportunity to redefine what it means to work on IR research? Can we be unabashedly discontent with imagining the future of IR based wholly on what new AI progress makes plausible, and instead reimagine our field as a place where knowledge, culture, and radical aspirations meet to demand of technology to make new futures possible? Alternatively, if we fail to articulate an aspiring vision for IR research, we risk as a field being reduced to just an application of AI.
LLMs are raising new sociotechnical concerns.
It is well-known that language models reproduce, and even amplify, harmful stereotypes and biases of moral import (Friedman and Nissenbaum, 1996) that are present in their training data (Bolukbasi et al., 2016; Caliskan et al., 2017; Gonen and Goldberg, 2019; Blodgett et al., 2020; Bender et al., 2021; Abid et al., 2021). One particular mitigation strategy involves using ML approaches that learn from human preferences, such as reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Ziegler et al., 2019), to align with “human values” (Kasirzadeh and Gabriel, 2023; Tamkin et al., 2021). While RLHF has been quite effective in constraining LLMs from producing certain types of offensive and harmful content, we must be wary of any framing of AI ethics, such as AI alignment (Russell et al., 2015; Gabriel, 2020; Gabriel and Ghazavi, 2021), that presupposes the existence of universal values but that assumption does not hold true in reality (Prabhakaran et al., 2022; Birhane and Cummins, 2019; Jobin et al., 2019; Png, 2022; Sambasivan et al., 2021). This is particularly concerning if we look at this in the context of power asymmetries that exist between powerful private corporations, who have outsized influence over what values these models are optimized for, and those who use these models or are represented in some fashion in the model outputs. This is further compounded by the lack of appropriate mechanisms for civil society to participate in and challenge these choices. Indeed, by placing these controls in the hands of the privileged few, we risk further concentration of power. The concerns of biases in what these models produce, and even biases in what context they refuse to generate (Urman and Makhortykh, 2023), and who gets to influence those decisions have serious implications for information access and society. Finally, the development of LLMs themselves may involve potential harms to authors (Davis, 2023; Lawler, 2023; Browne, 2023; Shetler, 2024; Milmo, 2024), crowdworkers (Gray and Suri, 2017; Ekbia and Nardi, 2017; Gray and Suri, 2019; Roberts, 2019; Jones-Imhotep, 2020; Roberts, 2021; Williams et al., 2022; Perrigo, 2023; Dzieza, 2023), and even the environment (Bender et al., 2021; Patterson et al., 2021; Bommasani et al., 2021; Wu et al., 2022a; Dodge et al., 2022; Patterson et al., 2022).
In a critical perspective, Shah and Bender (2022) recommend that IR research should focus on developing appropriate guardrails in anticipation of the social implications of these emerging technologies and not be constrained by a singular LLM-powered conversational search vision for IR. With this we agree, we should be excited by the new capabilities unlocked by recent progress in LLMs but must not limit our imaginations and aspirations by only what LLMs make plausible. We must also consider the broad sociotechnical implications of deploying these emerging technologies in the context of information access (Mitra et al., 2024), and their systemic consequences and risks.
Our relationships with adjacent communities are changing and so is how we do research.
Not long ago, many in the IR community would tout IR, and specifically web search, as a rare success story of real-world application of ML, AI, and NLP technologies. A sea change in these adjacent communities in the last decade have shifted this balance. Now, many see IR as just another NLP task, sometimes included in NLP benchmarks (e.g., HELM (Liang et al., 2022)) for evaluating ML and AI models. With retrieval-augmented LLMs, IR is auditioning for a new role, as a tool for AI models, curiously inverting the relationship between these technologies, where AI was one of many in the IR toolbox.
These communities are also undergoing significant changes in research culture, often influencing each other on the way. One particular trend in NLP, and the broader ML and AI communities, that has influenced IR, is leaderboard-driven research. Several NLP leaderboards (Bajaj et al., 2016; Goyal et al., 2017; Joshi et al., 2017; Rajpurkar et al., 2018; Wang et al., 2018; Talmor et al., 2018; Wang et al., 2019; Kwiatkowski et al., 2019; Yang et al., 2018; Liang et al., 2020) have been instrumental in encouraging progress on specific tasks. IR has a long history of focus on shared tasks and benchmarks, notably TREC (Voorhees et al., 2005) that has been a venue for developing new tasks and benchmarks, as well as building research communities with shared interests around them. What differentiates IR benchmarking in venues such as TREC from NLP leaderboards is that the former is framed not as a competition, but as a coopetition. In a competition, the goal of the participant is to outperform others, while in a coopetition the participants share a collective goal to develop a better understanding of both the task and the models in the spirit of scientific enquiry. In the words of IR researcher Ian Soboroff: [^9] https://twitter.com/ian_soboroff/status/1426901262369439751 “The datasets were not built to be solved. They were built as tools to understand the problem and the systems we build to ‘solve’ them.”; or as Voorhees (2021) put it: “Coopetition is defined as competitors cooperating for the common good… While competition can give one a bigger piece of the pie, cooperation makes the whole pie bigger.” By emphasizing the goals of community development and understanding of the tasks and the models, these evaluation effort try to promote scientific enquiry over sportive competition. Even the MS MARCO ranking task that initially started as a competition later reframed itself as a coopetitive evaluation effort (Craswell et al., 2021; Lin et al., 2022).
As a community we should cautiously embrace insights and trends from our neigboring fields. However, we should not let IR be minimized to just an evaluation task which undermines the critical responsibility that IR researchers owe to the broader society. Similarly, while leaderboards and competitions may be effective in creating excitement and increasing participation in certain tasks, we must be mindful of the implications of potentially a large section of the IR community being driven predominantly by these practices. When the goal is to win, then scientific inquiry takes a back seat, and the ones with the most compute and data resources take the metaphorical steering wheel. It risks, what Gausen et al. (2023) call, albeit in a different context, shifting from praxisi.e., “reflection and action directed at the structures to be transformed” (Freire, 2020)—to proxies, i.e., optimizing towards proxy quantitative measures of outcomes. Actions in this context may refer to any research activity, including but not limited to: formalization, design, experimentation, publishing, artefact creation, open sourcing, and community building; and examples of proxies include state-of-the-art (SOTA) performance on benchmarks and leaderboard rankings that do not translate to better scientific understanding or positive impact on people. [^10] This is a case of Goodhart’s law (Chrystal and Mizen, 2001; Goodhart, 1975; Hoskin, 1996; Thomas and Uminsky, 2022) whereby improvements on benchmarks and corresponding metrics do not translate to progress on the problem the benchmark was created for, as has been argued for example by Hsia et al. (2023).
Yet another relationship that we must critically examine is the one between industry and academia. Whittaker (2021) point out that the concentration of data and compute resources, two key ingredients in recent advances in AI, in the hands of few large tech corporations is giving these same institutions tremendous power to shape academic research agenda. Big tech also shapes academic research agendas in various other ways, including academic engagements and employments. In IR, the MS MARCO dataset (Bajaj et al., 2016) and leaderboard (Craswell et al., 2021; Lin et al., 2021, 2022), and the TREC Deep Learning track (Craswell et al., 2020), that has been broadly adopted for benchmarking deep ranking models were produced and is currently maintained by industry researchers. Indeed, the organizers behind these efforts themselves recognized (Craswell et al., 2021; Lin et al., 2022) the critical responsibility that comes with defining critical research tasks for the community—effectively playing “the Pied Piper guiding a significant section of the community down specific lanes of research”—and recommend all benchmark developers to engage in open and inclusive discussions with the rest of the community to critically examine their impact. While academia-industry collaboration is critically important for the field to ground our research in real large-scale systems and see our research outputs materialize into real impact on system users, we must also resist the homogenization of our research agendas towards a singular world view put forth by Big Tech capitalism.
The world is changing and so is our relationship to that world.
Our world at large is experiencing a confluence of many simultaneous, and mutually reinforcing, forces that are increasingly pushing us towards precarity, including but not limited to: increasing global wealth and income inequality (Chancel et al., 2022), rising global conflicts (Taylor, 2023; United Nations Meetings Coverage and Press Releases, ), pandemics (Scientist, 2021; Taylor, 2022; for Disease Control et al., 2022), and impending climate catastrophes (Parmesan et al., 2022; IPCC, 2013; Poynting and Rivault, 2024). At a moment when the world needs global solidarity built on trust and consensus, and informed citizenry with robust access to reliable information, online disinformation and misinformation are undermining both (Turrentine, 2022; Treen et al., 2020; Kata, 2010; Allcott and Gentzkow, 2017; Doubek, 2017; Beaumont et al., 2020; Zadrozny, 2024; Swenson and Fernando, 2023). While these complex global challenges require sophisticated and multifaceted response that spans across the political, legal, economic, and technological realms, one thing is for certain that information access research has a role to play. So, will we answer the call?
5 Conclusion
Despite good intentions.